OpenStudy (anonymous):

the base of a triangle is 4cm greater than the height. the area is 48cm^2 find teh base of the triangle

OpenStudy (anonymous):

A = (1/2)(b*h) 48 = (1/2)((4+h)h) 48 = (1/2)(4h+h^2) 96 = 4h + h^2 h^2 + 4h - 96 = 0 (h + 12)(h - 8) = 0 h = -12 or 8 since height cannot be a negative value, h = 8 cms therefore, b = 12 cms The height of the triangle is 8 centimeters. The base of the triangle is 12 centimeters.

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

same as ryleesmama